OpenStudy (anonymous):

Find the area of a circle with a diameter of 7 inches.

OpenStudy (divanshu):

find radius and then use pie*radius^2

OpenStudy (anonymous):

\[A = \pi r^2 = \pi (7/2)^2 = 49\pi / 4\]

OpenStudy (anonymous):

R = D/2, so R = D/2 :)

OpenStudy (anonymous):

38.48in^2?

OpenStudy (anonymous):

yes! ^_^